OpenStudy (anonymous):

Medal for the most creative answer... What is 0+0? And no, it does not have to be mathematical!!!

OpenStudy (mary.rojas):

0+0= LOVE,PEACE,HAPPINESS, AND FRENCH FRIES. :)

OpenStudy (anonymous):

hmm... that's pretty good.

OpenStudy (mary.rojas):

LOL

OpenStudy (anonymous):

How about mine? 0+0 = the essence of the multiverse and all of its being. You see, the zeros represent black holes that open into a space-time continuum. They lead you int alternate universes, and since each of these universes have infinite timelines and infinite black holes with infinite universes in those, therefore: 0+0=infinity.

OpenStudy (dumbcow):

creative but what evidence do you have of these alternate universes? :) 0+0 = 0 something can't come from nothing

OpenStudy (nory):

 Therefore, 0 + 0 = infinity.